OpenStudy (anonymous):

1.Suppose you pay $24 for a pair of shoes that has been discounted 20%. What is the original price of the shoes? Show how you identify what you are looking for, set up an equation, arrive at your answer, and check your work. Then clearly state your answer.

OpenStudy (anonymous):

$ 30

OpenStudy (anonymous):

$24 - 80% x - 100% x=24*100/80

OpenStudy (zarkon):

As my wife told me once...look at the price tag...they always show the original price in addition to any discounts :)

OpenStudy (anonymous):

you have paid $ 24 which is 20% less than original price let x be the original price x - 0.2x 24 0.8x = 24 x = 24/0.8 = 240/8 = $ 30